OpenStudy (anonymous):

anyone know how to differentiate [(a^2)(dx)]/ [x(sqrt(x^2 - a^2)]?

hartnn (hartnn):

there's a dx there..so, do u mean to ntegrate ?

hartnn (hartnn):

*integrate ?

OpenStudy (anonymous):

yup. i tried trigonometric substitution but i would somehow end up with the answer "a" only

OpenStudy (anonymous):

x = asecu dx = asecutanu sqrt(x^2 - a^2) = atanu therefor:

OpenStudy (anonymous):



hartnn (hartnn):

hey thats absolutely correct! but don't forget we haven't integrated yet!

hartnn (hartnn):

\(\huge \int a \: du=..?\)

OpenStudy (anonymous):

da?

hartnn (hartnn):

oh, and one error, \(dx = a \sec u\ tanu \: du\)

hartnn (hartnn):

not da, here the variable us 'u' , so 'du'

OpenStudy (anonymous):

oh. so it becomes au?

hartnn (hartnn):

correct :) au +c now substitute back for 'u' x=a sec u so, u=...?

OpenStudy (anonymous):

u = asec/x wait. does sec still have u or do i change it to x?

hartnn (hartnn):

u = asec/x ......no \(\large x = a \sec u \\ \large (x/a)=\sec u \\ \large u = \sec^{-1}(x/a)\) the inverse secant function!

hartnn (hartnn):

gotthis ?

OpenStudy (anonymous):

ahhhh... so it becomes a(sec^-1 [x/a]) right?

hartnn (hartnn):

correct! don't forget the +c part :)

OpenStudy (anonymous):

wait. does that mean i have to cancel a?

hartnn (hartnn):

you cannot cancel 'a' the other a is inside the inverse secant function

hartnn (hartnn):

a(sec^-1 [x/a]) +c will be the final answer

OpenStudy (anonymous):

ohh...got it! thank you :)

hartnn (hartnn):

welcome ^_^ ask if anymore doubts...

OpenStudy (anonymous):

btw. what if the other sie has y? a(sec^-1 [x/a]) = y (y was already integrated) a(sec^-1 [x/a]) = y + C (sec^-1 [x/a]) =( y + C)/a ?

hartnn (hartnn):

ohh...differential equations ? correct, go on... x/a = sec [(y+c)/a ] or x = a sec [(y+c)/a] ....

OpenStudy (anonymous):

ahhhh....so thats how it is. thanks har!! :D

hartnn (hartnn):

welcome ^_^